# TAROT: Task-Oriented Authorship Obfuscation Using Policy Optimization Methods

### v1

@@ Abstract
Authorship obfuscation aims to disguise the identity of an author within a text by altering the writing style, vocabulary, syntax, and other linguistic features associated with the text author. This alteration needs to balance privacy and utility. While strong obfuscation techniques can effectively hide the author’s identity, they often degrade the quality and usefulness of the text for its intended purpose. Conversely, maintaining high utility tends to provide insufficient privacy, making it easier for an adversary to de-anonymize the author. Thus, achieving an optimal trade-off between these two conflicting objectives is crucial. In this paper, we propose TAROT: Task-Oriented Authorship Obfuscation Using Policy Optimization, a new unsupervised authorship obfuscation method whose goal is to optimize the privacy-utility trade-off by regenerating the entire text considering its downstream utility. Our approach leverages policy optimization as a fine-tuning paradigm over small language models in order to rewrite texts by preserving author identity and downstream task utility. We show that our approach largely reduce the accuracy of attackers while preserving utility. We make our code and models publicly available. [^1] https://github.com/hornetsecurity/tarot
@@ 1 Introduction
Privacy is a critical consideration when developing and deploying trustworthy machine learning systems. With the expansion of large language models (LLM), large datasets are used for pretraining and fine-tuning models on a wide variety of tasks. With this comes a risk of sensitive or personal information being exposed or misused. Text anonymization is a technique that can address these privacy concerns. It involves the process of removing or obfuscating the author’s personal information from textual data. This helps protect the privacy of individuals whose data is used to train models, however, it needs to be performed in a way that still allows models to learn useful patterns and insights from texts.
@@
Currently, most work done on text anonymization focuses on redacting sensitive entities in a given document Lison et al. (2021). This is sufficient for texts where the only private aspects are entities, such as medical reports, court cases, or biographies. But it is inadequate for removing author’s writing style, or weak signals that can be used as hints for identification, which is for example the case for blog articles or emails. Redacting entities in text while keeping stylometric features linked to a specific individual would eventually result in a leak of information. Indeed, the writing style is a strong indicator of a person’s identity Mosteller and Wallace (1963). Prior work done on authorship attribution highlights the large quantity of information that can be extracted from seemingly anonymized texts and the ease of identification of authors, especially for long documents Fabien et al. (2020).
To solve this issue, authorship obfuscation (AO) aims to hide the author’s identity by replacing some part of the text associated with authorship indicators. Modifying the original text can impact its usability for specific tasks (i.e. utility), and therefore badly affects the downstream performances and text comprehension of machine learning models. The enforcement of privacy creates a trade-off between privacy and utility, where keeping the original text preserves the unchanged utility of the text, while not defending against attribution attacks. On the other side, obfuscating the entire text guarantees privacy, but leads to unusable text in practice.
Previous approaches design their obfuscation by maximizing the preserved text content. They limit the modifications to small and targeted edits in order to preserve text meaning and keep textual content as close as possible to the original. As a result, those approaches often lead to insufficient modification in the text, especially against realistic attack scenarios Zhai et al. (2022).
To address these limitations, we reframe the AO problem into an adversarial problem between two adversaries (e.g. machine learning models): one attacker model whose goal is to reveal the identity of a given author from written texts, and one utility model that aims to perform a given task using authors data. The goal is to provide a modified version of the original text such that the utility model can accurately perform its task while preventing the attacker from identifying the author, making the obfuscation task-oriented. The notion of task-oriented obfuscation/anonymization also takes its origin in the law. As stated by GDPR European Parliament and Council of the European Union (2016), the collection and process of personal information (including written texts) have to be specified for a given usage.
@@
We fine-tune a text simplification model for AO using a customized reward model. We design an unsupervised reward model for PO using 2 pretrained sentence embedding models. The utility reward, penalizes the fact that the General Text Embeddings Li et al. (2023) of the anonymized sentence is too far removed from that of the original sentence. The author rewards does the opposite on the embedding built by the Universal Authorship Representation model from Rivera-Soto et al. (2021). Our final models are trained in an open-world setting where the number of authors is not defined, same goes for the end utility for our model to work on a multi-task setting. We also provide experimentation on three different datasets, movie reviews, blog articles and scholar documents. We show that TAROT can be used on multiple datasets targeting different tasks while protecting authorship.
@@ 2 Related Work
Recently, ALISON Xing et al. (2024) employs a lightweight multilayer perceptron classifier using part-of-speech (POS) sequences to guide obfuscation, and leverages a BERT pre-trained language model to generate replacement sequences. By ranking and replacing important POS n-grams, ALISON obfuscates text uniformly, reducing classifier confidence.
Reinforcement Learning.
@@
With the expansion of Reinforcement Learning from Human Feedback (RLHF) as a LLM fine-tuning paradigm, RL techniques have been leveraged to improve language models with scalar metrics by optimizing rewards from (human) feedbacks. It has emerged as a prominent tool for tackling undesirable behaviors such as toxicity, social biases, and offensive language Ouyang et al. (2022). This is accomplished by implementing PO algorithms to optimize a language model (LM) by associating a reward with each generation, derived from a trained reward model.
@@ 3.1 Problem Formulation
Let $x_{\text{ori}}$ represent the original document authored by a specific author $a\in\mathcal{A}$. $\mathcal{A}$ denoting a predetermined set of authors. The objective of authorship obfuscation is to generate a new document, denoted as $x_{\text{obf}}$, which cannot be attributed to the original author $a$. To assess the effectiveness of obfuscation, we employ a classification model, denoted as $f_{priv}(\cdot)$ (i.e. an authorship attribution model), which has been trained to distinguish documents based on their respective authors within $\mathcal{A}$. The goal of authorship obfuscation is to design an obfuscation method $O(\cdot)$, such that $f_{priv}(O(x_{\text{ori}}))\neq f_{priv}(x_{\text{ori}}).$
In addition, a successful obfuscation algorithm would not only trick an attacker into predicting the wrong author, but also preserve the document utility for downstream usage. In this paper, instead of mainly measuring this utility change though various semantic or content preservation metrics (i.e. METEOR score, BERT score, …) we highlight the selection of a prior task $\mathcal{T}$ in order to evaluate obfuscation with respect to $\mathcal{T}$. We denote as $f_{\mathcal{T}}(\cdot)$ the classification model used for a utility task. An ideal $O(\cdot)$ would preserve the original label $f_{\mathcal{T}}(O(x_{\text{ori}}))=f_{\mathcal{T}}(x_{\text{ori}})$.
@@ 3.2 Framework Overview
3.3 Supervised fine-tuning
First, we use a fine-tuned LM to initiate our text generation task. We employ the Keep It Simple [^2] https://hf.co/philippelaban/keep_it_simple simplification model Laban et al. (2021) as an SFT baseline. This model is a fine-tuned version of GPT2-medium on the Newsela [^3] https://newsela.com/ dataset for text simplification. The utilization of a simplification model encourages a reduction in the amount of information conveyed by a sentence, thereby affording the opportunity to eliminate author-specific features. To our knowledge, this is the first time that a simplification model has been used for AO. Moreover, our framework is broadly compatible with any autoregressive LM, and can be adapted with larger architectures and other generation tasks.
@@ 4.1 Datasets
To evaluate our obfuscation models, we use three different datasets. (i) IMDb62555https://hf.co/datasets/tasksource/imdb62, is a subset of the IMDb Authorship Attribution dataset initially presented by Seroussi et al. (2014). It consists of 62 authors with 1,000 texts per author taken from IMDb movie reviews. The utility task associated with this dataset is the review sentiment. For this, we map the movie rating between 0 and 10 associated with each review to a sentiment between positive and negative. A positive review occurs when the review rating is strictly larger than 5. (ii) The Blog Authorship Corpus [^6] https://u.cs.biu.ac.il/~koppel/BlogCorpus.htm dataset Schler et al. (2006) consists of aggregated blog posts from 19,320 bloggers gathered from blogger.com. We pick the list of 13 topics present in the dataset as the utility task. (iii) The Extended-Brennan-Greenstadt [^7] https://hf.co/datasets/tasksource/Drexel-AMT dataset Brennan et al. (2012) is composed of short paragraphs about scholar subjects gathered from 42 different authors from Amazon Mechanical Turk. The utility task of this dataset is indicated by the “background” column, as a binary classification problem.
@@ 4.2 Reward Models
For the privacy reward, we use the Learning Universal Authorship Representations model (LUAR), from Rivera-Soto et al. (2021). LUAR’s goal is to transform a given text into a 512 dimensions embeddeding, such that representations of texts by the same author are closer, according to cosine similarity, than those by other authors.
@@
We obtain our final reward by summing the two previous rewards:
$R=R_{util}+R_{priv}$
All implementations details are listed in Appendix A.1.
@@ 4.3 Evaluation
As a baseline, we perform a naive text edition using synonyms. We use GPTZzz [^9] https://github.com/Declipsonator/GPTZzzs to process original texts, it employs a dictionary of synonyms to replace a given proportion of words with their counterparts. The goal of this baseline is to evaluate the attacker behavior when very small edits are made in the original text.
@@ 4.4 Realistic Attackers
So far, the considered privacy evaluation adopts the perspective of an attacker who does not have any knowledge about the obfuscation algorithm used. In a more realistic setting, the attacker can likely identify and reproduce the AO model, and perform more advanced attacks by creating adversarial threat models. Following Zhai et al. (2022), we also evaluate our obfuscation models against two enhanced authorship attribution attackers, better suited to simulate real world attack scenarios. We list the different attackers and their specific aspects:
  • Data Augmentation: The attacker is aware of which algorithm was used to perform AO. This knowledge is used to create a new attribution model combining texts from both the original source and obfuscated texts generated from the AO algorithm.
  • Adversarial: The attacker is also aware of the AO algorithm, but use the obfuscated training samples only as a training set for the authorship attribution classifier. Zhai et al. (2022) showed that this attack setting is the one having the higher performances on text edition obfuscations.
@@ 5 Results
Adversarial Attackers
Figure 2 highlights the accuracy of adversarial threat models on the IMDb-10 dataset, both attackers reduce the original attacker accuracy drop once retrained using obfuscated texts for the Synonyms approach and Alison. However, text generation methods (GPT-3.5, SFT, TAROT-PPO and TAROT-DPO) are resistant to adversarial threat models, and only GPT-3.5 and TAROT-DPO suffer from a performance decrease from the data augmentation attack. This encourages the path of generation methods as promising obfuscators. Note that this is the first obfuscation approach that is shown to be resistant to threat models. As Zhai et al. (2022) did not include generation models in their study of AO evaluation.
Utility Preservation after Retraining
Figure 2: Adversarial training accuracy results on IMDB-10 (lower is better). Generation models are resistant to adversarial training, compared to text edition methods.
@@
We show an obfuscation example in Table 3 for each method. The base Synonyms obfuscation results in awkward phrasing and less natural language, compromising readability. ALISON maintains coherence and clarity with slight formalization (“thoroughly enjoyed” instead of “loved”). GPT-3.5 significantly rephrases the text using sophisticated language. SFT simplifies and shortens the text, retaining clarity but reducing stylistic nuances. TAROT-PPO simplifies further, introducing some repetition, which makes the text less formal but still clear. TAROT-DPO alters the content more significantly, introducing new themes and repetition that can distract from the original meaning. The application of PO assists the text simplification SFT model in making additional modifications to the text. Although these changes in some cases alter the text’s meaning, they preserve its overall utility.
@@
As a complement, we perform an ablation study of each component of our reward model in Appendix D. It confirms the importance of using a combination of both privacy and utility rewards to learn this trade-off in obfuscation, especially for PPO.
@@ 6 Conclusion
We introduced a novel authorship obfuscation framework that focuses on optimizing the privacy-utility trade-off for a specific downstream data usage. We fine-tuned a text simplification model using two policy optimization algorithms to obfuscate the authorship of a given text, while preserving utility for multiple tasks. Our end-models are tuned using two sentence embedding rewards, one for content preservation and one for privacy, resulting in an unsupervised approach made for the open-world authorship setting. The results obtained help to improve the privacy from state-of-the-art AO methods, while preserving task utility. Additionally, we show that generated texts can be used to retrain utility classifier and increase their performances, while limiting the accuracy of more advanced attackers. This calls for more research to design robust evaluation benchmarks for obfuscation systems, to assess and catch failure cases that can map to different real-world scenarios.

### v2

@@ Abstract
Authorship obfuscation aims to disguise the identity of an author within a text by altering the writing style, vocabulary, syntax, and other linguistic features associated with the text author. This alteration needs to balance privacy and utility. While strong obfuscation techniques can effectively hide the author’s identity, they often degrade the quality and usefulness of the text for its intended purpose. Conversely, maintaining high utility tends to provide insufficient privacy, making it easier for an adversary to de-anonymize the author. Thus, achieving an optimal trade-off between these two conflicting objectives is crucial. In this paper, we propose TAROT: Task-Oriented Authorship Obfuscation Using Policy Optimization, a new unsupervised authorship obfuscation method whose goal is to optimize the privacy-utility trade-off by regenerating the entire text considering its downstream utility. Our approach leverages policy optimization as a fine-tuning paradigm over small language models in order to rewrite texts by preserving author identity and downstream task utility. We show that our approach largely reduces the accuracy of attackers while preserving utility. We make our code and models publicly available. [^1] https://github.com/hornetsecurity/tarot
@@ 1 Introduction
Text is a primary medium for storing user data, training machine learning models, and interacting with large language models (LLMs) during inference. However, it also poses significant privacy risks, as sensitive or personal information contained within text can be exposed or misused. Text anonymization is a vital technique to address these concerns by removing or obfuscating personal information. This process protects individual privacy while ensuring that machine learning models can still derive meaningful insights and patterns from anonymized data, preserving its utility.
@@
Currently, most work done on text anonymization focuses on redacting sensitive entities in a given document Lison et al. (2021). This is sufficient for texts where the only private aspects are named entities, such as medical reports, court cases, or biographies. But it is inadequate for removing the author’s writing style, or the weak signals that can be used as hints for identification, which is, for example, the case for blog articles or emails. Redacting entities in text while keeping stylometric features linked to a specific individual would eventually result in a leak of information. Indeed, the writing style is a strong indicator of a person’s identity Mosteller and Wallace (1963). Previous work on authorship attribution highlights the large amount of information that can be extracted from seemingly anonymized texts and the ease of identification of authors, especially for long documents Fabien et al. (2020).
To solve this issue, authorship obfuscation (AO) aims to hide the author’s identity by replacing some part of the text associated with authorship indicators. Modifying the original text can impact its usability for specific tasks (i.e. utility), and therefore badly affects the downstream performances and text comprehension of machine learning models. The enforcement of privacy creates a trade-off between privacy and utility, where keeping the original text preserves the unchanged utility of the text, while not defending against attribution attacks. On the other hand, obfuscating the entire text guarantees privacy, but leads to unusable text in practice. Previous approaches design their obfuscation by maximizing the preserved text content. They limit the modifications to small and targeted edits in order to preserve text meaning and keep textual content as close as possible to the original. While this strategy is necessary to maintain the exact content and ensure that we convey the exact same message (before publishing the text online for example), those approaches often lead to insufficient modification in the text, especially against realistic attack scenarios Zhai et al. (2022).
To address these limitations, we reframe the AO problem into an adversarial problem between two adversaries (e.g. machine learning models): one attacker model whose goal is to reveal the identity of a given author from written texts, and one utility model that aims to perform a given task using authors’ data. The goal is to provide a modified version of the original text such that the utility model can accurately perform its task while preventing the attacker from identifying the author, making the obfuscation task-oriented. This perspective is more angled towards data users who need to privately perform utility tasks on the data, where some degree of content alteration may be acceptable if it enhances privacy. The notion of task-oriented obfuscation/anonymization also takes its origin in the law. As stated by GDPR European Parliament and Council of the European Union (2016), the collection and processing of personal information (including written texts) must be specified for a given usage.
@@
We fine-tune a text simplification model for AO using a customized reward model. We design an unsupervised reward model for PO using two pretrained sentence embedding models. The utility reward penalizes the fact that the General Text Embeddings Li et al. (2023) of the anonymized sentence is too far removed from that of the original sentence. The author rewards does the opposite on the embedding built by the Universal Authorship Representation model from Rivera-Soto et al. (2021). Our final models are trained in an open-world setting where the number of authors is not defined, the same goes for the end utility for our model to work on a multi-task setting. We also provide experimentation on three different datasets, movie reviews, blog articles and scholar documents. We show that TAROT can be used on multiple datasets targeting different tasks while protecting authorship.
@@ 2 Related Work
Recently, ALISON Xing et al. (2024) employs a lightweight multilayer perceptron classifier using part-of-speech sequences to guide obfuscation, and leverages a BERT pre-trained language model to generate replacement sequences. By ranking and replacing important part-of-speech n-grams, ALISON obfuscates text uniformly, reducing classifier confidence.
Related studies share a common approach to evaluating privacy: they measure it through the performance of authorship attribution classifiers against obfuscated texts. Zhai et al. (2022) push forward this evaluation framework by introducing adversarial attackers that can resist obfuscation techniques. For measuring utility, the standard is to treat AO as a reference-less natural language generation problem, and to rely on standard metrics used for similar tasks such as machine translation and summarization Altakrori et al. (2022).
Reinforcement Learning
@@
With the development of Reinforcement Learning from Human Feedback (RLHF) as a LLM fine-tuning paradigm, RL techniques have been leveraged to improve language models with scalar metrics by optimizing rewards from (human) feedback. It has emerged as a prominent tool for tackling undesirable behaviors such as toxicity, social biases, and offensive language Ouyang et al. (2022). This is accomplished by implementing PO algorithms to optimize a language model (LM) by associating a reward with each generation, derived from a trained reward model.
@@
Private Synthetic Text Generation
Our work lies at the frontier between private text editing and synthetic text generation. Creating private synthetic data often relies on established frameworks such as differential privacy Dwork (2006). In contrast to these approaches, we focus on the implementation of a single text-to-text transformation specifically designed for authorship obfuscation, rather than on the generation of new textual data derived from potentially multiple sources Mattern et al. (2022a).
Differential privacy traditionally targets noise addition in documents to produce useful and private text representations Feyisetan et al. (2019); Fernandes et al. (2019). Applying differential privacy to document rewriting primarily serves to mitigate membership inference attacks, addressing a distinct threat model compared to the authorship attribution attacks targeted by our approach. While these techniques exhibit emergent capabilities for masking authorship signals Igamberdiev and Habernal (2023); Weggenmann et al. (2022); Utpala et al. (2023), they typically do so at a substantial cost to text utility, both at the task-level and the syntactic-level Mattern et al. (2022b). This approach introduces unnecessary noise to semantic content not relevant to authorship identification, often degrading the overall coherence and readability of the text. In contrast, our obfuscation methodology implements targeted modifications to stylometric features while maintaining the overall integrity of the source text.
@@ 3.1 Problem Formulation
Let $x_{\text{ori}}$ represent the original document authored by a specific author $a\in\mathcal{A}$. $\mathcal{A}$ denoting a predetermined set of authors. The objective of authorship obfuscation is to generate a new document, denoted as $x_{\text{obf}}$, which cannot be attributed to the original author $a$. To assess the effectiveness of obfuscation, we employ a classification model, denoted as $f_{attr}(\cdot)$ (i.e. an authorship attribution model), which has been trained to distinguish documents based on their respective authors within $\mathcal{A}$. The goal of authorship obfuscation is to design an obfuscation method $O(\cdot)$, such that $f_{attr}(O(x_{\text{ori}}))\neq f_{attr}(x_{\text{ori}}).$
In addition, a successful obfuscation algorithm would not only trick an attacker into predicting the wrong author, but also preserve the document utility for downstream usage. In this paper, instead of mainly measuring this utility change though various semantic or content preservation metrics (i.e. METEOR score, BERT score, etc.) we highlight the selection of a prior task $\mathcal{T}$ in order to evaluate obfuscation with respect to $\mathcal{T}$. We denote as $f_{\mathcal{T}}(\cdot)$ the classification model used for a utility task. An ideal $O(\cdot)$ would preserve the original label $f_{\mathcal{T}}(O(x_{\text{ori}}))=f_{\mathcal{T}}(x_{\text{ori}})$.
@@ 3.2 Framework Overview
3.3 SFT Initialization
First, we use a fine-tuned LM to initiate our text generation task. We employ the Keep It Simple [^2] https://hf.co/philippelaban/keep_it_simple simplification model Laban et al. (2021) as an SFT baseline. This model is a fine-tuned version of GPT2-medium on the Newsela [^3] https://newsela.com/ dataset for text simplification. The utilization of a simplification model encourages a reduction in the amount of information conveyed by a sentence, thereby affording the opportunity to eliminate author-specific features [^4] Our preliminary experiments revealed that using a simplification model outperformed comparable models of similar size for copy, paraphrasing, back-translation, and summarization, delivering superior privacy and utility.. To our knowledge, this is the first time that a simplification model has been used for AO. Moreover, our framework is broadly compatible with any autoregressive LM, and can be adapted with larger architectures and other generation tasks.
@@ 4.1 Datasets
To evaluate our obfuscation models, we use three different datasets. (i) IMDb62666https://hf.co/datasets/tasksource/imdb62, is a subset of the IMDb Authorship Attribution dataset initially presented by Seroussi et al. (2014). It consists of 62 authors with 1,000 texts per author taken from IMDb movie reviews. The utility task associated with this dataset is the review sentiment. For this, we map the movie rating between 0 and 10 associated with each review to a sentiment between positive and negative. A positive review occurs when the review rating is strictly larger than 5. (ii) The Blog Authorship Corpus [^7] https://u.cs.biu.ac.il/~koppel/BlogCorpus.htm dataset Schler et al. (2006) consists of aggregated blog posts from 19,320 bloggers gathered from blogger.com. We pick the list of 13 topics present in the dataset as the utility task. (iii) The Extended-Brennan-Greenstadt [^8] https://hf.co/datasets/tasksource/Drexel-AMT dataset Brennan et al. (2012) is composed of short paragraphs about scholar subjects gathered from 42 different authors from Amazon Mechanical Turk. The utility task of this dataset is indicated by the “background” column, as a binary classification problem.
@@ 4.2 Reward Models
For the privacy reward, we use the Learning Universal Authorship Representations model (LUAR), from Rivera-Soto et al. (2021). LUAR’s goal is to transform a given text into a 512 dimensions embedding, such that representations of texts by the same author are closer, according to cosine similarity, than those by other authors.
@@
We obtain our final reward by summing the two previous rewards $R=R_{util}+R_{priv}$. All implementation details are listed in Appendix A.1.
@@ 4.3 Evaluation
As a baseline, we perform a naive text edition using synonyms. We use GPTZzzs [^10] https://github.com/Declipsonator/GPTZzzs to process original texts, it employs a dictionary of synonyms to replace a given proportion of words with their counterparts. The goal of this baseline is to evaluate the attacker behavior when very small edits are made in the original text.
@@ 4.4 Realistic Attackers
So far, the considered privacy evaluation adopts the perspective of an attacker who does not have any knowledge about the obfuscation algorithm used. In a more realistic setting, the attacker can likely identify and reproduce the AO model, and perform more advanced attacks by creating adversarial threat models. Following Zhai et al. (2022), we also evaluate our obfuscation models against two enhanced authorship attribution attackers, better suited to simulate real-world attack scenarios. We list the different attackers and their specific aspects based on adversarial training:
  • Mix of original and obfuscated texts: The attacker knows which AO algorithm was used and leverages this knowledge to create a new attribution model. This model is trained on a combination of original source texts and obfuscated texts generated by the known AO algorithm. We use a 50/50 distribution between original and obfuscated data to train this attacker.
  • Only obfuscated texts: While the attacker is also aware of the AO algorithm, they train their authorship attribution classifier exclusively on the obfuscated samples. Zhai et al. (2022) demonstrated that this attack setting achieves the highest performance against text edition obfuscations.
@@ 5 Results
Adversarial Attackers
Figure 2 highlights the accuracy of adversarial threat models on the IMDb-10 dataset. This attack strategy is effective against text edition approaches (Synonyms and ALISON) as shown by the accuracy gain compared to the base attack only trained on original texts. However, text generation methods (GPT-3.5, SFT, TAROT-PPO and TAROT-DPO) show resistance to adversarial threat models, and only GPT-3.5 and TAROT-DPO are susceptible to the attacker trained on a mix of original and obfuscated texts. This encourages the path of generation methods as promising obfuscators. Note that this is the first obfuscation approach that is shown to be resistant to threat models. [^11] Zhai et al. (2022) did not include generation models in their study of AO evaluation.
Figure 2: Authorship adversarial training accuracy results on IMDB-10 (lower is better). Generation models are resistant to adversarial training, compared to text edition methods.
Utility Preservation After Retraining
@@
We show an obfuscation example in Table 3 for each method. The base Synonyms obfuscation results in awkward phrasing and less natural language, compromising readability. ALISON maintains coherence and clarity with slight formalization (“thoroughly enjoyed” instead of “loved”). GPT-3.5 significantly rephrases the text using sophisticated language. SFT simplifies and shortens the text, retaining clarity but reducing stylistic nuances. TAROT-PPO simplifies further, introducing some repetition, which makes the text less formal but still clear. TAROT-DPO alters the content more significantly, introducing new themes and repetition that can distract from the original meaning. The application of PO assists the text simplification SFT model in making additional modifications to the text. Although these changes in some cases alter the text’s meaning, they preserve its overall utility. Appendix F provides more obfuscation examples from proposed and baseline methods.
@@
As a complement, we perform an ablation study of each component of our reward model in Appendix D. It confirms the importance of using a combination of both privacy and utility rewards to learn this trade-off for obfuscation, especially for PPO.
@@ 6 Conclusion
We introduced a novel authorship obfuscation framework that focuses on optimizing the privacy-utility trade-off for a specific downstream data usage. We fine-tuned a text simplification model using two policy optimization algorithms to obfuscate the authorship of a given text, while preserving utility for multiple tasks. Our end-models are tuned using two sentence embedding rewards, one for content preservation and one for privacy, resulting in an unsupervised approach made for the open-world authorship setting. The results obtained help to improve the privacy from state-of-the-art AO methods, while preserving task utility. Our findings suggest that editing approaches are not suitable for privacy, especially against realistic attack settings. Additionally, we show that generated texts can be used to retrain utility classifiers and increase their performances, while limiting the accuracy of more advanced attackers. Ultimately, the performance of obfuscation methods largely varies depending on the downstream task choice, as does the resulting privacy-utility trade-off, highlighting the importance of selecting an appropriate model based on the specific requirements of the intended application. This calls for more research to design robust evaluation benchmarks for obfuscation systems, to assess and catch failure cases that can map to different real-world scenarios.